A balloon is filled with 3.5 L of water at 24 C. The balloon is placed outside on a hot day, 32 C, causing the volume to increase to ____ L. (only type your numerical answer and include 2 decimal places.


1
Expert's answer
2021-05-17T04:09:20-0400

Solution:

Water is a liquid at 24°C

The thermal volume coefficient (β) of water is 0.000207°C-1 from 20°C to 30 °C and 0.000303°C-1 from 30°C to 40°C.

(https://www.engineeringtoolbox.com/volumetric-temperature-expansion-d_315.html)


V = Vo + VoβΔT


At 30°C,

ΔT = (30 - 24)°C = 6°C

V = (3.5 L) + (3.5 L × 0.000207°C-1 × 6°C) = 3.504 L


At 32°C,

ΔT = (32 - 30)°C = 2°C

V = (3.504 L) + (3.504 L × 0.000303°C-1 × 2°C) = 3.506 L = 3.51 L

V = 3.51 L


Answer: 3.51